Draw line/s for each item;

3. Find the equation of the line parallel to 5x + 6y – 10 = 0 at a distance 61 from (-3,7).

4. Find the equation of the line parallel to 3x + 4y = 20 and at a distance of 5 units from this

line.


1
Expert's answer
2021-10-18T03:45:22-0400

4.

Given equation line, 3x−4y+2=0∴4y=3x+2y=(4/3)x+1/2​Slope of given line=3/4​As slope of parallel lines are equal∴ Slope of line parallel to given line =3/4​Equations of line passes through (−2,3) with slope (3/4​) is(y−3)=3/4​(x+2)4y−12=3x+63x−4y+18=0 be the equation of line parallel to 3x−4y+2=0 and passing through (−2,3)